# Neural Concept Verifier: Scaling Prover-Verifier Games via Concept Encodings

### v2

@@ 4.2 Evaluations
Lastly, to assess whether NCV can mitigate shortcut learning in image classification, we train models on different versions of CLEVR-Hans7 with varying ratios of clean samples (i.e., without shortcut) in the training and validation sets. We then measure validation accuracy with shortcuts and test accuracy on a held-out, clean data split. This setup allows us to track both predictive performance and robustness to shortcut learning. Table  2 reports results for three model types: a linear CBM, a non-linear CBM, and our instantiation of NCV using NCB as concept extractor. We observe that while NCV achieves the highest test accuracy among all models in the 0% clean data setting, it still exhibits a sizeable validation-test gap, indicating a strong influence of the underlying shortcuts. Notably, NCV shows a smaller gap than the non-linear CBM, suggesting that the inductive bias of the PVG-based training objective encourages a smaller influence of the shortcut, even without access to clean training data. As we progressively increase the amount of clean samples (from 1% to 20%), test accuracy and test-validation gap improves across all models. However, NCV consistently achieves the highest test accuracy in every setting, and its validation–test gap decreases more rapidly than for either CBM variant. This trend indicates that NCV is not only better at leveraging clean supervision when available, but is also more robust to shortcut learning. A similar trend is observed for CLEVR-Hans3 (cf. Table  6), where NCV closes the generalization gap even more efficiently than all baselines. Importantly, the linear CBM approach maintains a much larger gap than both non-linear approaches, with NCV providing additional soundness guarantees over the non-linear CBM. Together, these results demonstrate that concept-level Prover–Verifier Games in NCV encourage models to base decisions on more robust, task-relevant features. As a result, NCV is more resilient to shortcut learning, inparticular with access to minimal amounts of clean data.
Table 2: NCV shows increased robustness to shortcuts on CLEVR-Hans7. We report validation accuracy on a set with shortcuts, and test accuracy on a set without shortcuts, as well as the validation–test gap (lower is better) across models trained with varying amounts of clean data.
Ratio Clean (Samples) Model Val Acc (w/ shortcut) Test Acc (w/o shortcut) Val-Test Gap (↓)
[…]
NCV 98.38± $\mathbf{98.38}\mbox{\scriptsize$\pm\,0.18$}$ 92.23± $\mathbf{92.23}\mbox{\scriptsize$\pm\,0.67$}$ $6.15$
[…]
NCV 98.19± $\mathbf{98.19}\mbox{\scriptsize$\pm\,0.24$}$ 94.68± $\mathbf{94.68}\mbox{\scriptsize$\pm\,0.64$}$ $\mathbf{3.51}$
[…]
NCV 98.47± $\mathbf{98.47}\mbox{\scriptsize$\pm\,0.24$}$ 96.24± $\mathbf{96.24}\mbox{\scriptsize$\pm\,0.71$}$ $\mathbf{2.23}$
[…]
NCV 98.63± $\mathbf{98.63}\mbox{\scriptsize$\pm\,0.13$}$ 97.74± $\mathbf{97.74}\mbox{\scriptsize$\pm\,0.28$}$ $\mathbf{0.89}$
@@ 5 Discussion
Overall, our results show that shifting Prover–Verifier Games (PVGs) to the concept level yields a powerful and scalable framework for verifiable, interpretable classification. By operating on symbolic concept embeddings, NCV avoids the computational cost of per-sample inference in pixel space, yet matches or surpasses pixel-based baselines in both completeness and soundness. It reduces the performance gap typical of Concept Bottleneck Models (CBMs), achieving parity with opaque models on synthetic tasks and even surpassing them on natural images. Additionally, concept-level outputs offer concise, human-readable explanations. Finally, NCV exhibits a resilience to spurious correlations, generalizing from confounded training splits and closing the generalization gap with minimal injections of clean data.
That said, NCV has several limitations. Its effectiveness depends on the quality of the underlying concept extractor: noisy or entangled concept vocabularies can reduce both accuracy and human understandability. The increased training complexity introduced by the three-agent PVG setup also results in greater computational cost and training instability, e.g., compared to linear CBMs. Moreover, when using pretrained models like CLIP for concept discovery, NCV inherits their biases and inconsistencies to some extent (Birhane et al.,, 2021; Gehman et al.,, 2020; Bhalla et al.,, 2024). Finally, recent work (Debole et al.,, 2025) shows that such concept spaces can diverge from expert semantics, even when yielding strong downstream performance.
@@ 6 Conclusion
In this work, we have introduced the Neural Concept Verifier (NCV), a unified framework that brings together Prover–Verifier Games and concept-level representations for interpretable classification at scale. Through extensive experiments on CLEVR-Hans, CIFAR-100, and ImageNet-1k, we have shown that NCV achieves high completeness and soundness, reduces the interpretability–accuracy gap of concept bottleneck models, delivers detailed concept-based explanations, and effectively mitigates shortcut learning. Thus, NCV paves the way for deploying trustworthy and transparent models in domains where both predictive performance and verifiability are essential.
Future work should explore how concept encodings can be integrated into alternative PVG-style setups, where structured representations may improve performance or reduce communication overhead. It is also promising to investigate applications beyond vision, such as natural language processing and structured data, where interpretable verification may be equally valuable. Finally, while existing information-theoretic guarantees, such as those introduced by Wäldchen et al., (2024), focus on binary classification under specific assumptions, extending such guarantees to high-dimensional, multi-class settings remains an important open direction for formal interpretability at scale.
@@ Acknowledgements
This research was partially supported by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy – The Berlin Mathematics Research Center MATH+ (EXC-2046/1, project ID: 390685689). This work was further supported by the ”ML2MT” project from the Volkswagen Stiftung, and the ”The Adaptive Mind” project from the Hessian Ministry of Science and Arts (HMWK). It has further benefited from the HMWK projects ”The Third Wave of Artificial Intelligence - 3AI”, and Hessian.AI, and the research collaboration between TU Darmstadt and Aleph Alpha Research through Lab1141.

### v3

@@ 4.2 Evaluations
Lastly, to assess whether NCV can mitigate shortcut learning in image classification, we train models on different versions of CLEVR-Hans3 and CLEVR-Hans7 with varying ratios of clean samples (i.e., without shortcut) in the training and validation sets. We then measure validation accuracy with shortcuts and test accuracy on a held-out, clean data split. This setup allows us to track both predictive performance and robustness to shortcut learning. Tab.  2 reports results for three model types: a linear CBM, a nonlinear CBM, and our instantiation of NCV using NCB as concept extractor. We observe that while NCV achieves the highest test accuracy among all models in the 0% clean data setting, it still exhibits a sizeable validation-test gap, indicating a strong influence of the underlying shortcuts. As the amount of clean samples is progressively increased, test accuracy and test-validation gap improves across all models. However, NCV consistently achieves the highest test accuracy in every setting, and its validation–test gap decreases more rapidly than for either CBM variant. This trend indicates that NCV is not only better at leveraging clean supervision when available, but is also more robust to shortcut learning. Together, these results demonstrate that concept-level prover–verifier interaction in NCV encourage models to rely on more robust, task-relevant features, making NCV more resilient to shortcut learning, even with limited amounts of clean data.
@@ 5 Discussion
Overall, our results show that shifting Prover–Verifier Games (PVGs) to the concept level yields a powerful and scalable framework for verifiable, interpretable classification. By operating on symbolic concept embeddings, NCV avoids the computational cost of per-sample inference in pixel space, yet matches or surpasses pixel-based baselines in both completeness and soundness. It reduces the performance gap typical of Concept Bottleneck Models (CBMs), achieving parity with opaque models on synthetic tasks and even surpassing them on natural images. Concept-level outputs offer concise, human-readable explanations. Finally, NCV exhibits higher resilience to spurious correlations, generalizing from confounded training splits and closing the generalization gap with minimal available clean data.
That said, NCV has several limitations. Its effectiveness depends on the quality of the underlying concept extractor: noisy or entangled concept spaces can reduce both accuracy and human understandability. The increased training complexity introduced by the three-agent PVG setup also results in greater computational cost and training instability, e.g., compared to linear CBMs. Moreover, when using pretrained models like CLIP for concept discovery, NCV inherits their biases and inconsistencies to some extent (Birhane et al., 2021; Gehman et al., 2020; Bhalla et al., 2024). Finally, recent work (Debole et al., 2025) shows that such concept spaces can diverge from expert semantics, even when yielding strong downstream performance.
@@ 6 Conclusion
In this work, we have introduced the Neural Concept Verifier (NCV), a unified framework that combines Prover–Verifier Games (PVGs) and concept‐level representations for interpretable classification at scale. Through extensive experiments on CLEVR-Hans, CIFAR-100, ImageNet-1k, and COCOLogic, we have shown that NCV scales well to high-dimensional datasets by achieving high completeness and soundness, reduces the interpretability–accuracy gap of concept bottleneck models, delivers detailed concept-based explanations, and effectively mitigates shortcut learning. Thus, NCV provides a promising pathway for deploying trustworthy and transparent models in domains where both predictive performance and verifiabile explanations are essential.
Future work should explore how concept encodings can be integrated into alternative PVG-style setups, where structured representations may improve performance or reduce communication overhead. It is also promising to investigate applications such as natural language processing and structured data, where interpretable verification may be equally valuable. Finally, a compelling direction for future research lies in the optimization level; specifically, developing more stable schemes for training Merlin and Morgana end-to-end on discrete, binarized masks. Advancing these discrete concept selection techniques could further strengthen both the framework and the provers themselves.
@@ Acknowledgements
This research was partially supported by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy – The Berlin Mathematics Research Center MATH+ (EXC-2046/1, EXC-2046/2, project ID: 390685689). This work was further supported by the ”ML2MT” project from the Volkswagen Stiftung, and the ”The Adaptive Mind” project from the Hessian Ministry of Science and Arts (HMWK). It has further benefited from the HMWK projects ”The Third Wave of Artificial Intelligence - 3AI”, and Hessian.AI, the EU-funded “TANGO” project (EU Horizon 2023, GA No 57100431), from early stages of the Cluster of Excellence "Reasonable AI" funded by the German Research Foundation (DFG) under Germany’s Excellence Strategy— EXC-3057 and the research collaboration between TU Darmstadt and Aleph Alpha Research through Lab1141. It has also benefited from the DFG project GRK 2853 "Neuroexplicit Models of Language, Vision, and Action" (project number 471607914).